# On the hierarchical Bayesian modelling of frequency response functions

### v1

@@ 6.1 Model development
Prior to fitting the FRF model, random Gaussian noise was generated, with a magnitude equal to 5% of the absolute peak value of the FRF from Blade 2. (Noise was added to the training data to avoid over-fitting small deviations/lightly participating modes within the band of interest.) The noise was then added to each FRF, and 100, 100, 7, and 20 training points were randomly selected from the FRFs belonging to Blades 1 to 4, respectively. The intent was that the two data-rich FRFs (Blades 1 and 2) would lend statistical support to the sparser domains (Blades 3 and 4), thereby reducing uncertainty compared to an approach without pooling. Recall the grouping of the peaks from the FRFs shown in Figures 2a and 2b. Both data-rich FRFs belonged to one of the groupings; likewise, the two data-poor FRFs belonged to the other grouping. This represents a challenging situation where the model can be heavily biased towards the available training data. The hierarchical/partial-pooling modelling structure allows the target distributions to be informed by the data (for very data-poor domains, this will of course be limited), but with the help of the full population dataset (including the data-rich domains). The Stan HMC sampler was run using four chains, with a target average proposal acceptance probability rate of 0.99 (this parameter controls the target acceptance rate of the NUTS algorithm, and setting it to a value close to 1 can reduce the number of false-positive divergences M. Betancourt (2017)). The sampler was run for 10000 samples per chain (and an additional 5000 warm-up samples were discarded per chain to diminish the influence of the starting values A. Gelman et al. (2013)), for each parameter. In addition, models were run for each set of blade data independently (no pooling), for comparison with the partial-pooling model.
@@ 6.2 Model results
FRFs for each blade were computed via Eq. (13), from the samples of the modal parameters. Total variance was estimated by adding the standard deviation of the FRFs to the expectation of the noise variance. Posterior predictive mean and 3-sigma deviation for the partial pooling and independent models are plotted in Figures 6a to 7b, respectively.
@@
Indeed, Figures 6a to 7b show that the sparser FRFs improved significantly for the combined model, compared to the independent models. With only 7 data points, Blade 3, shown in Figure 7a in green, was missing most of the information necessary to characterise the two modes. The independent model for Blade 3 relied heavily on the user-specified prior, while the partial-pooling model borrowed information from the other three blades to inform the shared latent model. Similar (albeit less severe), results were seen for Blade 4, shown in Figure 7b in purple. With 20 data points, the FRF for Blade 4 was less sparse than that for Blade 3, but was still somewhat lacking near the resonance peaks, especially for the first mode. Figure 7b shows a significant reduction in variance with the partial-pooling approach. The improvements in the data-rich FRFs, Blades 1 (blue) and 2 (red) were not as apparent, as expected. Variance reduction from the combined-inference (partial-pooling) approach can also be visualised by plotting the marginal distributions of the hyperparameters. For the population-level variables, marginal distributions were approximated by sampling the expectation and variance for each variable (or shape parameters, for damping), and then drawing from a distribution (normal for natural frequency and residue, and beta for damping) with the same statistics as the parameter draws. Kernel density estimation (KDE) was then used to approximate marginal distributions using the population-level draws. Marginal distributions for the domain-level variables were approximated via KDE of the samples. For the natural frequencies, the population- and domain-level distributions for each mode are shown in Figures 8a and 8b. For the (shared) residues, which were learnt at the population-level, the parent- and lower-level distributions for each mode are shown in Figures 9a and 9b. For damping, the population- and domain-level distributions for each mode are shown in Figures 8c and 8d.
@@
Figure 8: Marginal distributions for natural frequency, (a) mode 1 and (b) mode 2; and for damping, (c) mode 1 and (d) mode 2. Population- and domain-level distributions are shown at the top and bottom, respectively. For the partial-pooling model, population-level distributions are shaded and grey, and domain-level distributions are shaded and blue (Blade 1), red (Blade 2), green (Blade 3), and purple (Blade 4). For the independent (no-pooling) models, population- and domain-level distributions are shown as dashed lines, using the same colour scheme as the partial-pooling model.
@@
Figure 9: Marginal distributions for residue, (a) mode 1 and (b) mode 2. Parent- and lower-level distributions are shown at the top and bottom, respectively. For the partial-pooling model, parent- and lower-level distributions are shaded and grey. For the independent (no-pooling) models, parent- and lower-level distributions are shown as dashed lines, where Blade 1 is shown in blue, Blade 2 in red, Blade 3 in green, and Blade 4 in purple.
@@
Figures 8a to 9b show that when the domains were able to share information via common higher-level distributions, the variance was reduced compared to a no-pooling approach. In Figure 8a, for the first natural frequency, the population-level distribution of the partial-pooling model was taller and with lower variance than the independent models, which was likely because the data-rich domains dominated the available data for that mode, compared to the data-poor domains (forcing the distribution towards the natural frequencies of the data-rich domains). This data sharing was also evident at the domain-level, where the KDEs of the data-poor domains aligned closely with those from the data-rich domains, for the partial-pooling models. The no-pooling KDEs of the data-poor FRFs differed significantly, and were primarily informed by the priors. In contrast, Figure 8b shows that for the second natural frequency, the population-level distribution for the partial-pooling model was flatter than those for the independent models. This was likely to have occurred because more data were available for the second peak from Blade 4, which was from the second grouping (widening the distribution to accommodate the greater differences in natural frequency). Again, this conclusion is supported at the domain-level, where the KDE for Blade 4 is very similar for the partial-pooling and independent models (although there is still a small variance reduction for the partial-pooling model). For Blade 3, which was quite data-poor at the second mode as well as the first, the KDE in Figure 8b shows that the natural frequency was estimated to occur in between those of the other, more data-rich blades. Again, for Blade 3, the no-pooling KDE differed significantly from the others, as this parameter was primarily informed by the priors, because of a lack of data at the second mode. Figures 8c and 8d show that the population-level distributions for damping did not vary significantly among the partial-pooling and independent models, likely because the priors were fairly strong and also accurate to the data. At the domain-level, KDEs for the data-rich FRFs showed high similarity for the different pooling models, while KDEs for the data-poor FRFs tended towards the priors. Similar results were seen for the residues, in Figures 9a and 9b, except that the residues were shared among the domains for the partial-pooling model, so there was only one lower-level distribution for each mode. Close alignment of the domain-level distributions, of the data-rich domains, suggests that sharing the residue among the different blades was a suitable assumption. The results presented in this case demonstrate the development of a population form, where a combined hierarchical FRF model is learnt for a group of nominally-identical helicopter blades. Two domains (Blades 3 and 4) had limited data, and were especially sparse near the resonance peaks. Such a situation could occur if an insufficient frequency resolution was selected during the data acquisition process. By borrowing data from data-rich domains within the population, variance was reduced compared to an independent modelling approach.
@@ 7.1 Identification of appropriate functional relationships between temperature and the modal parameters
Figure 10a shows that a second-order polynomial appears to be an appropriate fit for natural frequency in the temperature range of interest. From Figure 10b, a linear fit was considered most appropriate for damping. Although there is likely a (weakly) nonlinear relationship between temperature and damping for this dataset, a linear assumption provided good results for the experiments used here. A higher-order Taylor series expansion for the temperature-damping relationship would require many coefficients, which would significantly increase the number of hyperparameters in the model. Further studies, including material-properties evaluation, may help elucidate the nature of this relationship specific to the helicopter blades, which may allow for better inferences from increased physics-based knowledge.
@@ 7.2 Model development and results
As with the first case, the population model for the second case included four latent functions (i.e., $K=4$), as four temperature-varied FRFs provided training data, i.e., those at -10, -5, 10, and ${}^{\circ}$. Likewise, only the real part of the FRF was fitted for simplicity. The real components of the FRFs were modelled probabilistically with an assumed Gaussian distribution, using the same accelerance FRF estimation as a likelihood function, as in Eqs. 13 and 18. The additive noise variance $\sigma^{2}_{\text{H}}$ of the FRFs was assumed shared among each of the domains (again, the same data acquisition system and sensors were used among the different tests), and the modal residue was assumed constant with respect to temperature. Shared, higher-level distributions were placed over natural frequency, $\boldsymbol{\omega}_{nat}=\{\omega_{nat}^{k}\}_{k=1}^{4}$, damping, $\boldsymbol{\zeta}=\{\zeta_{k}\}_{k=1}^{4}$, and residue, $A$, to allow information sharing among the domains,
@@
Note that hyperpriors for $\boldsymbol{\omega}_{nat}$ are shown in rad/s. Also note that for this preliminary model, normal distributions were assumed for all parameters (with truncated normal distributions assumed for variance, damping, and natural frequency). The residue, $A$, was assumed constant over all temperatures, with distributions,
@@
The shared noise variance $\sigma^{2}_{\text{H}}$ was expressed via,
$\sigma^{2}_{\text{H}}\sim\mathcal{TN}\left(0.3,1^{2}\right)\$ (28)
@@
Prior to fitting the FRF model, random Gaussian noise was generated, with a magnitude equal to 5% of the absolute peak value of the FRFs. The noise was then added to each FRF, and 100 training points were randomly selected from each of the FRFs captured at temperatures -10, -5, 10, and ${}^{\circ}$. The intent was to use a sub-set of the temperature-varied FRFs, to learn the population-level coefficients necessary to describe the relationships between temperature and the modal parameters, so that these population-level variables can be used to make predictions at ‘unseen’ temperatures (of course, these predictions were validated using experiments that did not contribute to the training data). As with the previous case, the Stan HMC sampler was run using four chains (and with a target average proposal acceptance probability rate of 0.99), for 10000 samples per chain (with an additional 5000 warm-up samples per chain, which were discarded), for each parameter. The expectation for the domain and population-level parameters were obtained from the samples. The expectation of the domain-level natural frequencies and damping, and shared residue and noise variance, were used to compute the FRFs via Eq. (13), as shown in Figure 7.2. The expectation of the population-level variables which were used to extrapolate to other temperatures are shown in Table 1. (Because the shared residue and noise variance were assumed to be temperature-invariant, they were also used to extrapolate to other temperatures).
@@ 7.3 Extrapolation to other temperatures
After training the model, predictions were made at ‘unseen’ temperatures, not used in training, using the expectation of population-level variables (shown in Table 1), and compared to measured FRFs to evaluate model accuracy. All shared variables (i.e., residue, polynomial coefficients, and noise variance), were assumed equal to the expectation of the variables computed using the model. Natural frequency and damping were estimated for all the measured temperatures, $T=\{T_{i}\}_{i=1}^{11}$, including the four used in training, via the relations,
$\begin{split}&\{\omega_{nat}^{i}=\text{E}[{\mu}_{\omega_{nat}}]+\text{E}[{a_{1}}]T_{i}+\text{E}[{a_{2}}]T_{i}^{2}\}_{i=1}^{11}\\ &\{\zeta_{i}=\text{E}[{\mu}_{\zeta}]+\text{E}[{b}]T_{i}\}_{i=1}^{11}\\ \end{split}$ (29)
The extrapolated FRFs are shown in Figure 7.3, ranging from -20 to ${}^{\circ}$ in increments of ${}^{\circ}$, with decreasing temperature from left to right, as natural frequency increases with decreasing temperature. Model accuracy was evaluated by calculating the NMSE via Eq. (15) for each measured FRF with added noise, as shown in Table 2. Note that the NMSE computed for the prediction at temperatures -10, -5, 10, and ${}^{\circ}$, which correspond to the data used in model training, are included in the table for completeness. However, the specific training points were excluded from the calculation. To further visualise the relationships between temperature and the modal parameters, the functions were computed for each sample. The mean and variance of the functions were plotted with the parameters estimated via independent modelling (in this case, these results were considered to be the best approximation of the ground truth, as each FRF was data-rich). These plots are shown in Figures 12a and 12b.
@@
In Figure 7.3, the extrapolated FRFs are shown as solid blue lines, with a shaded blue region indicating the variance bounds. The measured FRFs used to train the model, without added noise, are shown as solid red lines, and the measured FRFs used to test the model at ‘unseen’ temperatures are shown as dashed black lines. From the figure, it is clear that excellent agreement was achieved between the extrapolated and measured FRFs at the temperatures used to train the model, as expected. Likewise, the FRFs at temperatures between those used to train the model (e.g., at 0, 15, and ${}^{\circ}$) show excellent agreement. Notably, the FRFs at colder temperatures, which were further away from the training data, still show good agreement. The accuracy of the fit is further shown in Table 2, as all NMSE values were less than 5%. Figures 12a and 12b further show that the extrapolated parameters accurately represented the measured data, as the measured parameters largely fell within the variance bounds of the approximations. Figure 12a shows that for natural frequency, parameter variance increased as the inferred parameters varied further from the training data (i.e., at temperatures further from those used to train the model), as expected. Figure 12b shows that for damping, the relatively consistent variance was the result of the linear assumption for the temperature-damping relationship, and the relatively high deviation of the measured data from the linear fit. Further model development could involve the incorporation of more physics-based knowledge, such as, forcing the relationship between natural frequency and temperature to be monotonically decreasing, or further investigation into the possibly nonlinear relationship between damping and temperature.

### v2

@@ 6.1 Model development
The models were trained using data with added noise, to avoid over-fitting small deviations/lightly-participating modes within the band of interest. This was accomplished by adding zero-mean Gaussian noise, with variance equal to 10% of the absolute maximum amplitude of the data from Blade 1, to the 20 time-data blocks prior to computing FRFs. Then, the time data were subsampled such that there were two data-rich domains (Blades 1 and 2), and two data-poor domains (Blades 3 and 4). (Note that the data-rich groups were subsampled as well. This procedure was adopted to reduce the dimensionality of the training data, although there were enough training points to characterise each mode within the band of interest for these blades.) FRFs were then computed and averaged, and truncated to the desired bandwidth. This procedure resulted in 281, 207, 11, and 31 training points for Blades 1, 2, 3, and 4, respectively.
The intent was that the two data-rich FRFs would lend statistical support to the sparser domains, thereby reducing uncertainty compared to an approach without pooling. (Note that partial pooling is also applicable to other situations, such as having only a single data-rich member, as briefly explored in A.) Recall the grouping of the peaks from the FRFs shown in Figures 2a and 2b. Both data-rich FRFs belonged to one of the groupings; likewise, the two data-poor FRFs belonged to the other grouping. This represents a challenging situation where the model can be heavily biased towards the available training data. The hierarchical/partial-pooling modelling structure allows the target distributions to be informed by the data (for very data-poor domains, this will of course be limited), but with the help of the full population dataset (including the data-rich domains). In addition, the subsampling and added noise resulted in some outliers in the data, creating a challenging inference problem. The Stan HMC sampler was run using four chains, with a target average proposal acceptance probability rate of 0.8 (this parameter controls the target acceptance rate of the NUTS algorithm, and setting it to a value close to 1 can reduce the number of false-positive divergences M. Betancourt (2017)). The sampler was run for 10000 samples per chain (and an additional 5000 warm-up samples were discarded per chain to diminish the influence of the starting values A. Gelman et al. (2013)), for each parameter. Models were also run for each set of blade data independently (no pooling), for comparison with the partial-pooling model. Posterior predictive checks were performed to ensure stationarity and proper mixing of the Markov chains. As computed by Stan, the maximum split $R_{hat}$ diagnostic, and the minimum effective sample size ($N_{Eff}$), considering all parameters/hyperparameters, are shown in Table 1, for all models. Note that split $R_{hat}$ values that do not exceed 1.01 suggest proper mixing of the chains A. Vehtari et al. (2021). Likewise, total $N_{Eff}$ values greater than 400 are considered suitable for stable estimates and proper interpretation of split $R_{hat}$, for a setup comprised of four parallel chains, as employed in this work A. Vehtari et al. (2021). These criteria were met for all models developed.
Table 1: Posterior predictive diagnostics for Case 1.
Partial Pooling
No Pooling
Metric
All Blades
Blade 1
Blade 2
Blade 3
Blade 4
min $N_{Eff}$
3085
518
7583
1290
1251
max $R_{hat}$
1
1
1
1
1
@@ 6.2 Model results
Eq. (13) was used to compute FRFs from the posterior MCMC samples of the modal parameters, for each blade. FRFs with predicted variance were generated by sampling from a Student’s t-distribution with four degrees-of-freedom, mean equal to the computed FRFs, and posterior predictive measurement noise, ($\sigma^{2}$), for each sample. Total variance was estimated by taking the standard deviation of the FRFs. Posterior predictive mean and 3-sigma deviation for the partial-pooling and independent models are plotted in Figures 6a to 7b, respectively.
@@
Indeed, Figures 6a to 7b show that the sparser FRFs improved significantly for the combined model, compared to the independent models. With only 11 data points, Blade 3, shown in Figure 7a in green, was missing most of the information necessary to characterise the two modes. The independent model for Blade 3 relied heavily on the user-specified prior, while the partial-pooling model borrowed information from the other three blades to inform the shared latent model. Similar (albeit less severe), results were seen for Blade 4, shown in Figure 7b in purple. With 31 data points, the FRF for Blade 4 was less sparse than that for Blade 3, but was still somewhat lacking near the resonance peaks, especially for the first mode. Figure 7b shows a significant reduction in variance with the partial-pooling approach. The improvements in the data-rich FRFs, Blades 1 (blue) and 2 (red) were not as apparent, as expected. Variance reduction from the combined-inference (partial-pooling) approach can also be visualised by plotting the marginal distributions of the parameters/hyperparameters. For the population-level variables, marginal distributions were approximated by taking the posterior MCMC samples for expectation and variance for each variable (or shape parameters, for damping), and then drawing from a distribution (normal for natural frequency and residue, and beta for damping) with the same statistics as the parameter draws. Kernel density estimation (KDE) was then used to approximate marginal distributions using the population-level draws. (Note it is known what distribution a given sample from the posterior will take, e.g., expectation and variance are learnt assuming a priori a normally-distributed mode shape. However, the shape of the posterior distribution is itself unknown, i.e., the distribution over these distributions. Because the posterior is itself a random variable, KDE can be used as an approximation for the marginal distributions, using the posterior MCMC samples.) Marginal distributions for the domain-level variables were also approximated via KDE of the posterior MCMC samples. For the natural frequencies, the population- and domain-level distributions for each mode are shown in Figures 8a and 8b. For damping, the population- and domain-level distributions for each mode are shown in Figures 8c and 8d. For the (shared) residues, which were learnt at the population-level, the parent- and lower-level distributions for each mode are shown in Figures 9a and 9b.
@@
Figure 8: Marginal distributions for natural frequency, (a) mode 1 and (b) mode 2; and for damping, (c) mode 1 and (d) mode 2. Population- and domain-level distributions are shown at the top and bottom, respectively. For the partial-pooling model, posterior distributions are shaded. For the independent (no-pooling) models, posterior distributions are shown as dashed lines. Priors are shown as solid black lines.
@@
Figure 9: Marginal distributions for residue, (a) mode 1 and (b) mode 2. Parent- and lower-level distributions are shown at the top and bottom, respectively. For the partial-pooling model, posterior distributions are shaded. For the independent (no-pooling) models, posterior distributions are shown as dashed lines. Priors are shown as solid black lines.
@@
Figures 8a to 9b show that when the domains were able to share information via common higher-level distributions, the variance was reduced compared to a no-pooling approach. In these figures, the population-level distributions are shown at the top, while domain-level distributions are shown on the bottom of the plots. In Figure 8a, for the first natural frequency, the population-level distribution of the partial-pooling model was closely aligned with those of the independent models for the data-rich domains, and was taller and with lower variance than those of the independent models for the data-poor domains. This observation was likely because the data-rich domains dominated the available data for that mode, compared to the data-poor domains (forcing the distribution towards the natural frequencies of the data-rich domains). This data sharing was also evident at the domain-level, where the KDEs of the data-poor domains aligned closely with those from the data-rich domains, for the partial-pooling models. Also, in Figure 8a, it can be seen that the no-pooling KDEs of the data-poor FRFs were primarily informed by the priors. Indeed, the population-level distribution for the sparsest domain, Blade 3, was nearly indistinguishable from the prior, indicating that the data had little influence on the posterior (this makes sense, as there were very little data to characterise this mode). In contrast, Figure 8b shows that for the second natural frequency, the population-level distribution for the partial-pooling model was flatter than those for the independent models. This was likely to have occurred because more data were available for the second peak from Blade 4, which was from the second grouping (widening the distribution to accommodate the greater differences in natural frequency). Again, this conclusion is supported at the domain-level, where the KDE for Blade 4 is very similar for the partial-pooling and independent models (although there is still a small variance reduction for the partial-pooling model). For Blade 3, which was quite data-poor at the second mode as well as the first, the KDE in Figure 8b shows that the natural frequency was estimated to occur in between those of the other, more data-rich blades. Again, for Blade 3, the no-pooling KDE differed significantly from the others, as this parameter was primarily informed by the priors, because of a lack of data at the second mode. For both modes, it is clear that the independent models struggled to fit the data-poor FRFs, as evidenced by bimodal distributions for natural frequency. Figures 8c and 8d show that the population-level distributions for damping did not vary significantly among the partial-pooling and independent models (with only a slight deviation from the priors for the partial-pooling and data-rich models), likely because the priors were fairly strong and also accurate to the data. At the domain-level, KDEs for the data-rich FRFs showed high similarity for the different pooling models, while KDEs for the data-poor FRFs tended towards the priors. Similar results were seen for the residues, in Figures 9a and 9b, except that the residues were shared among the domains for the partial-pooling model, so there was only one lower-level distribution for each mode. Close alignment of the domain-level distributions, of the data-rich domains, suggests that sharing the residue among the different blades was a suitable assumption. The results presented in this case demonstrate the development of a population form, where a combined hierarchical FRF model is learnt for a group of nominally-identical helicopter blades. Two domains (Blades 3 and 4) had limited data, and were especially sparse near the resonance peaks. Such a situation could occur if an insufficient sampling rate was selected during the data acquisition process. By borrowing data from data-rich domains within the population, variance was reduced compared to an independent modelling approach.
@@ 7.1 Identification of appropriate functional relationships between temperature and the modal parameters
Figure 10a shows that a second-order polynomial appears to be an appropriate fit for natural frequency in the temperature range of interest. From Figure 10b, a linear fit was considered most appropriate for damping. Although there may be a (weakly) nonlinear relationship between temperature and damping for this dataset, a linear assumption provided good results for the experiments used here. A higher-order Taylor series expansion for the temperature-damping relationship would require many coefficients, which would significantly increase the number of hyperparameters in the model. Further studies, including material-properties evaluation, may help elucidate the nature of this relationship specific to the helicopter blades, which may allow for better inferences from increased physics-based knowledge.
@@ 7.2 Model development and results
As with the first case, the population model for the second case included four latent functions (i.e., $K=4$), as four temperature-varied FRFs provided training data, i.e., those at -10, -5, 10, and ${}^{\circ}$. In addition, the models developed were trained using data with added noise, to avoid over-fitting small deviations/lightly participating modes within the band of interest. This was accomplished by adding zero-mean Gaussian noise, with variance equal to 10% of the absolute maximum amplitude of the FRF obtained at ${}^{\circ}$, to the 20 time-data blocks prior to computing FRFs. Then, the time data were subsampled such that all four FRFs used in training were data rich (In other words, each FRF used in training had enough training points to fully characterise the mode in this band. Subsampling in this case was performed not to demonstrate shrinkage towards the data-rich groups, but to reduce the dimensionality of the training data). FRFs were then computed and averaged, and truncated to the desired bandwidth. This procedure resulted in 122 training points from each of the four temperature-varied FRFs. Also, once again, only the real part of the FRF was fitted for simplicity. (Again, given that the imaginary part could be modelled in the same fashion, or inferred from the real part, focussing on the real part only was deemed sufficient for demonstrating the techniques presented herein.) The real components of the FRFs were modelled probabilistically with an assumed Student’s t-distribution, using the same accelerance FRF estimation as a likelihood function, as in Eqs. 13 and 18. (Note that as with the previous case, $\nu$ was assumed to be four. This assumption was made for consistency with the previous case, although it could be learnt as a model parameter if desired.) The additive noise variance, $\sigma^{2}$, of the FRFs was assumed shared among each of the domains (again, the same data acquisition system and sensors were used among the different tests, and the same Gaussian noise were added), and the modal residue was assumed constant with respect to temperature. Shared, higher-level distributions were placed over natural frequency, $\boldsymbol{\omega}_{nat}=\{\omega_{nat}^{k}\}_{k=1}^{4}$, damping, $\boldsymbol{\zeta}=\{\zeta_{k}\}_{k=1}^{4}$, and residue, $A$, to allow information sharing among the domains,
@@
Note that hyperpriors for $\boldsymbol{\omega}_{nat}$ are shown in rad/s. Also note that for this preliminary model, normal distributions were assumed for all parameters (with truncated normal distributions assumed for variance, damping, and natural frequency). The only exception is the noise variance, which was modelled as a half-Cauchy, as with the previous case. The residue, $A$, was assumed constant over all temperatures, with distributions,
@@
The shared noise variance, $\sigma^{2}$, was sampled from a half-Cauchy parent distribution, with higher-level scale parameter, $\gamma^{2}$, sampled from a truncated normal distribution,
$\begin{split}&\sigma^{2}\sim\mathcal{C}\left(0,\gamma^{2}\right)\\ &\gamma^{2}\sim\mathcal{TN}\left(0.01,0.05^{2}\right)\\ \end{split}$ (28)
@@
As stated previously, FRFs captured at temperatures -10, -5, 10, and ${}^{\circ}$ provided the training data for the model. The intent was to use a sub-set of the temperature-varied FRFs, to learn the population-level coefficients necessary to describe the relationships between temperature and the modal parameters, so that these population-level variables can be used to make predictions at ‘unseen’ temperatures (of course, these predictions were validated using experiments that did not contribute to the training data). As with the previous case, the Stan HMC sampler was run using four chains (and with a target average proposal acceptance probability rate of 0.8), for 10000 samples per chain (with an additional 5000 warm-up samples per chain, which were discarded), for each parameter. Posterior predictive checks were performed to ensure stationarity and proper mixing of the Markov chains. As computed by Stan, the maximum split $R_{hat}$ diagnostic, and the minimum effective sample size ($N_{Eff}$), considering all parameters/hyperparameters, are shown in Table 2, for all models (including those used for ‘ground truth’ natural frequency and damping estimates). Note that split $R_{hat}$ values did not exceed 1.01, and $N_{Eff}$ was greater than 400, suggesting proper mixing of the Markov chains, per current guidelines A. Vehtari et al. (2021).
Table 2: Posterior predictive diagnostics for Case 2.
Metric
Partial-Pooling Model
‘Ground-Truth’ Models
min $N_{Eff}$
15000
640
max $R_{hat}$
1
1
Eq. (13) was used to compute FRFs from the posterior MCMC samples of the modal parameters, for each temperature-varied FRF used in training. FRFs with predicted variance were generated by sampling from a Student’s t-distribution with four degrees-of-freedom, mean equal to the computed FRFs, and posterior predictive measurement noise, $\sigma^{2}$, for each MCMC sample. Total variance was estimated by taking the standard deviation of the FRFs. Posterior predictive means and 3-sigma deviation for the FRFs are shown in Figure 7.2. The expectations for the population-level parameters were obtained from the samples, including the polynomial coefficients that define the relationships between temperature and the modal parameters. The expectations are shown in Table 3. The procedure for extrapolating beyond the FRFs used in training is described in Section 7.3.
@@ 7.3 Extrapolation to other temperatures
After training the model, predictions were made at ‘unseen’ temperatures, i.e., those not used in training. These predictions were made by computing natural frequency and damping using posterior MCMC samples, for all the measured temperatures, $T=\{T_{i}\}_{i=1}^{11}$, including the four used in training, via the relations,
$\begin{split}&\{\{\omega_{nat}^{i,n}={\mu}_{\omega_{nat,n}}+{a_{1,n}}T_{i}+{a_{2,n}}T_{i}^{2}\}_{i=1}^{11}\}_{n=1}^{40000}\\ &\{\{\zeta_{i,n}={\mu}_{\zeta,n}+{b_{n}}T_{i}\}_{i=1}^{11}\}_{n=1}^{40000}\\ \end{split}$ (29)
Natural frequency and damping were computed for each each MCMC sample, for each temperature. These values were then used to compute FRFs via Eq. (13). Variance was estimated by taking the standard deviation of the FRFs. The extrapolated FRF means are plotted in Figure 7.3, along with 3-sigma deviation. The extrapolated FRFs shown in Figure 7.3 range from -20 to ${}^{\circ}$ in increments of ${}^{\circ}$, with decreasing temperature from left to right, as natural frequency increases with decreasing temperature. Model accuracy was evaluated by calculating the NMSE via Eq. (15) for each measured FRF without added noise, compared to the mean extrapolated FRFs, as shown in Table 4. Note that the NMSE computed for the prediction at temperatures -10, -5, 10, and ${}^{\circ}$, which correspond to the data used in model training, are included in the table for completeness. However, the specific training points were excluded from the calculation. To further visualise the relationships between temperature and the modal parameters, the functions were plotted for each sample. The mean and variance of the functions were plotted with the parameters estimated via independent modelling (in this case, these results were considered to be the best approximation of the ground truth, as each FRF was data-rich). These plots are shown in Figures 12a and 12b.
@@
In Figure 7.3, the extrapolated FRFs are shown as solid blue lines, with a shaded blue region indicating the variance bounds. The measured FRFs used to train the model, without added noise, are shown as solid red lines, and the measured FRFs used to test the model at ‘unseen’ temperatures are shown as dashed black lines. From the figure, it is clear that excellent agreement was achieved between the extrapolated and measured FRFs at the temperatures used to train the model, as expected. Likewise, the FRFs at temperatures between those used to train the model (e.g., at 0, 15, and ${}^{\circ}$) show excellent agreement. Notably, the FRFs at colder temperatures, which were further away from the training data, still show good agreement. The accuracy of the fit is further shown in Table 4, as all NMSE values were less than 5%. Figures 12a and 12b further show that the extrapolated parameters accurately represented the measured data, as the measured parameters largely fell within the variance bounds of the approximations. Figure 12a shows that for natural frequency, parameter variance increased as the inferred parameters varied further from the training data (i.e., at temperatures further from those used to train the model), as expected. Figure 12b shows that for damping, the relatively consistent variance was the result of the linear assumption for the temperature-damping relationship, and the relatively high deviation of the measured data from the linear fit. Further model development could involve the incorporation of more physics-based knowledge, such as, forcing the relationship between natural frequency and temperature to be monotonically decreasing, or further investigation into the possibly nonlinear relationship between damping and temperature. In addition, it may be possible to augment the model by imposing a sparsity-inducing prior on the weights of the function describing the relationship. This sparsity-inducing prior would allow the Bayesian inference to perform system identification with automatic relevance detection, allowing incorporation of additional terms, polynomial or otherwise.